OpenStudy (anonymous):

Which does NOT pass through the origin? y=-2x^2 2x=square root of y 2y=negative absolute value of x y=-2x^2 + 2

ganeshie8 (ganeshie8):

test the options one by one

ganeshie8 (ganeshie8):

put x = 0, and see wat u get for y

ganeshie8 (ganeshie8):

take first option :- y = -2x^2 put x = 0 y = -2(0)^2 = 0 so, this passes through origin !

ganeshie8 (ganeshie8):

similarly test other options and see ?

OpenStudy (anonymous):

the last one would give me y=2, so that is not an origin. Am I correct?

ganeshie8 (ganeshie8):

perfect !

ganeshie8 (ganeshie8):

good job !!

OpenStudy (anonymous):

Thanks!!

ganeshie8 (ganeshie8):

np :)